4b) Calculate the half-width of a spectral line of wavelength  = 550 nm when the

temperature of the gas is 6  105 K. Assume H atoms to be emitters.


1
Expert's answer
2021-05-19T20:40:05-0400

The line shape is Gaussian; the full width at half maximum intensity in Å,



"\Delta\lambda_{1/2}^D = (7.16\times10^{-7})\,\lambda(T/M)^{1/2}\quad (1)"

where T is the temperature of the emitters in K, and M the atomic weight in atomic mass units (amu)

In our case, λ = 550 nm, T=6 × 105 K, M=1 amu

Using (1) we got:

"\Delta\lambda_{1/2}^D = (7.16\times10^{-7})\,\times550×10^{-9}(\frac{6×10^5}{1})^{1/2}"

"=3\times10^{-10}"


"\Delta\lambda_{1/2}^D = 3Å"

Answer:

"3\times10^{-10}"